### v1

@@ III-F Discussion
Each of the three evaluation methods possesses distinct advantages and drawbacks. Discriminative evaluation is often the method of choice for a range of standard benchmarks [6, 24]. This approach is intuitive, and the construction of evaluation prompts is straightforward. Selective evaluation resembles discriminative evaluation but is marginally less demanding because it includes a reference option for contrast. In both discriminative and selective evaluations, certain models might be suspected of conjecturing answers from few shots due to inadequate reasoning skills, which can undermine the reliability of the outcomes. Consequently, the use of explainable prompting becomes essential. Generative evaluation most closely mirrors real-world applications. However, the generated content is unrestricted, which poses challenges for even the most dependable reference-based evaluation techniques. Therefore, employing a combination of metrics simultaneously, including lexical evaluation based on token coverage and semantic evaluation based on textual similarity, is imperative.
@@
The foundational capabilities required of LLMs can be arrayed on a spectrum from simple to complex: generative, selective, and discriminative evaluation. Generative evaluation entails the direct invocation of parameters for continuation, bypassing the need for an extensive grasp of instructions, which suits models with minimal fine-tuning. Selective evaluation necessitates a degree of inferential reasoning but offers comparative choices, rendering the level of difficulty moderate. Conversely, discriminative evaluation demands the precise retrieval of factual information, thereby increasing the challenge.
Moreover, various evaluations cater to different application contexts. Should the objective be to solely improve the model’s capacity for reliable continuation, generative evaluation would suffice. In the training of a dependable chatbot, selective and discriminative evaluations prove suitable. When aiming to train a reward model, selective evaluation is beneficial, offering evaluation for positive and negative instances. If the goal is to enhance the model’s ability to recall and apply knowledge, discriminative evaluation emerges as the demanding option.
IV Related Works
IV-A Large Language Models
Language models are pivotal in computer science, evolving from statistical language models, to neural language models, to pre-trained language models (PLMs), and now to the current generation of LLMs. The advent of models such as ChatGPT has seen contemporary LLMs exhibit new capabilities in handling complex tasks. These models can manage few-shot tasks via in-context learning and tackle mixed tasks by following instructions [1]. LLMs can be classified according to two dimensions. The first dimension concerns the openness of the model weights. For example, open-source models include Meta’s LLaMA [17], Tsinghua University’s GLM [12], and Alibaba’s Qwen [14], while closed-source models feature OpenAI’s GPT [20], Baidu’s ERNIE Bot [31], and Anthropic’s Claude  [^5] https://www.anthropic.com/index/introducing-claude, among others. The second dimension differentiates between the use of a PLM or a supervised fine-tuned (SFT) model for specific inferences. A PLM is a language model trained on extensive unlabeled textual data to discern underlying patterns, structures, and semantic knowledge within the corpus. Conversely, an SFT model involves further training a PLM with labeled datasets tailored to a specific task, with the goal of improving performance in that area. Many open-source models, including LLaMA, GLM, and Qwen, have made their PLM weights publicly available. For SFT models, users can access the chat variants of open-source models or the API services provided by closed-source models. In our research, we focus primarily on evaluating closed-source GPT series models and open-source Chinese chat models.
TABLE VII: Hallucination Evaluation Benchmarks Sorted by Name
@@ IV-A Large Language Models
Benchmark (Released Year) Generation Method Annotation Metric Granularity Lang.
ChineseFactEval’23 [32] Manual Manual Acc Sentence CN
CSK-PN’23 [33] Direct: Common KGs No Need Acc Word EN
FACTOR’23 [10] CHG: Wiki, News Auto FACTOR Acc Sentence EN
FActScore’23 [9] CHG: Wiki No Need FActScore by Human Short Sentence EN
HaLoCheck’23 [34] CHG No Need HaLoCheck, selfcheckGPT Sentence EN
@@
FactualityPrompts’22 [35] Direct: Wiki Auto NE Error, Entailment Document, Sentence EN
HADES’22 [7] CHG: Wiki Manual Acc, G-Mean, BSS, AUC, etc. Word EN
@@
HalluQA’23 [24] CHG, Manual: TruthfulQA, Wiki Manual, Auto Non-hallucination Rate Sentence CN
HaluEval’23 [6] CHG: Alpaca, HotpotQA, etc. Manual, Auto Acc Document EN
HILT’23 [2] CHG: NYT, Politifact Manual HVI Word EN
@@
KoLA-KC’23 [36] Direct: Wiki, evolving dataset Auto BLEU, ROUGE Document EN
Med-HALT’23 [37] Direct: MedMCQA, PubMed, etc. No Need Acc, Pointwise Score All EN
PHD’23 [8] CHG: Wiki Manual F1, Acc, Prec, Reca Document EN
SelfAware’23 [38] CHG: Quora, HowStuffWorks Manual F1, Acc Sentence EN
STSN’23 [39] UHG Manual Acc, Prec, Reca Sentence, Concept EN
TruthfulQA’22 [28] Manual Manual Acc by Human or GPT-judge Sentence EN
@@
$\mathbb{UHG}$Eval (Ours) UHG: Xinhua News Auto, Manual Acc, kwPrec, BERTScore, etc. Sentence, Keyword CN
@@
XSum Hallu’20 [40] UHG: XSum Manual ROUGE, BERTScore, Acc, etc. Word, Document EN
@@
  • Note: Generation Method column provides the approach, and the base dataset if used. In this column, CHG refers to constrained hallucination generation, UHG refers to unconstrained hallucination generation, Manual indicates manually constructed, and Direct implies utilizing the base dataset without the need for generation. In the Annotation column, Auto denotes automatic machine annotation. In the Metric column, Acc, Prec, and Reca respectively indicate Accuracy, Precision, and Recall. In the Lang. column, CN and EN respectively stand for Chinese and English.
IV-B Hallucinations in LLM
Despite remarkable advancements in LLMs, they continue to encounter challenges, with hallucination being one of the most notable. Hallucination in language models refers to generating content that strays from factual accuracy, leading to unreliable outputs. Hallucinations occur when the generated content is not aligned with user input, deviates from the model’s previous outputs, or is at odds with established real-world knowledge [5]. Specific examples include inaccuracies in age, currency, scores, and other numerical values; citing fictional statements; inventing non-existent characters; and muddling timelines by merging events from different periods [2]. Regarding the causes of hallucinations, several factors can be responsible [5]. One contributing factor is the use of inaccurate or incomplete training data. During training, LLMs fine-tune their parameters with vast quantities of text data. However, this data may be flawed, harboring errors, inaccuracies, or gaps in information. Another factor involves inconsistencies in contextual information. While LLMs typically consider previously generated context when producing content, challenges in managing long-term dependencies or understanding complex contexts can result in inconsistencies. Additionally, hallucinations can arise from lacking or erroneous world knowledge. Although LLMs gain considerable world knowledge via training data, they may be deficient in specific domain knowledge or misinterpret certain facts, leading to hallucinations. Furthermore, model limitations, including generation strategies and alignment methods, can also play a role in hallucinations during content creation.
IV-C Hallucination Evaluation Benchmarks
To more effectively tackle the issue of hallucinations, constructing evaluation benchmarks is essential. In this context, numerous outstanding contributions have surfaced. This section reviews existing contributions regarding the development of benchmark datasets, their characteristics, and the particular methodologies for evaluation. Basic information about these benchmarks is presented in Table VII.
@@ IV-C Hallucination Evaluation Benchmarks
Benchmark dataset construction. Dataset construction usually involves three steps. Firstly, real-world texts for hallucination generation are collected, and most benchmarks directly use existing datasets, such as Wiki [10], Alpaca [6], PubMed [37], Quora [38] and so on. Secondly, hallucinations are generated usually by LLMs such as GPT3.5-Turbo, and most works uses constrained hallucination generation (CHG) paradigm [10, 9, 34, 6, 2, 8, 38]. STSN [39] and XSum Hallu [40] are the only two benchmarks that use UHG as we do. Thirdly, it is not certain that the content generated by the LLMs actually contains hallucinations, and often requires annotation, which is mostly done by human involvement. There are also works using automatic machine labeling [10, 35, 24, 6, 36]. These are the basic methods for constructing datasets, but there are also some other paradigms, such as constructing the dataset purely using manual labor, e.g. ChineseFactEval [32], HADES [7], TruthfulQA [28], etc.
@@
Benchmark dataset characteristics. Regarding the granularity of hallucinations labeled in the datasets, most studies assess hallucinations at the sentence and document levels, while a few examine them at the word (or keyword, concept) level. With respect to the topic, the majority of datasets cover the general domain, while some benchmarks target specific domains; for instance, HaLoCheck [34] focuses on the NBA, Med-HALT [37] on medicine, and our $\mathbb{UHG}$Eval on news. Concerning language, most evaluation datasets are in English. To our knowledge, the only two Chinese benchmarks, ChineseFactEval [32] and HalluQA [24], contain only 125 and 450 questions, respectively. Given the notably limited size of these datasets, our work significantly enhances the pool of data available for Chinese hallucination evaluation.
@@
Evaluation scheme. Existing works use a variety of ways to measure hallucinations. However, due to cost and time constraints, building automatic metrics for evaluation is still dominant, and a small proportion of works use human evaluation [9, 28, 40]. In terms of specific evaluation metrics, most works adopt common classification metrics, e.g., F1, accuracy, precision, recall. some other works construct their own calculation methods, e.g., FACTOR [10], FActScore [9], HaLoCheck [34], HVI [2], etc. However, the above metrics are rule-based and can only evaluate the ability of LLMs to classify hallucinations, but not the ability of LLMs to generate content without hallucinations. Thus, some benchmarks explore even further in generative evaluation. For example, KoLA [36] evaluates knowledge creation (KC) using BLEU and ROUGE to measure the degree of overlap between the output and the reference, TruthfulQA [28] evaluates hallucinations using a specially trained classifier, GPT-judge, and FactualityPrompts [35] simultaneously employs a hallucinated named entity error based on n-gram coverage and a semantic-based entailment ratio.
V Conclusion
@@
LLMs are experiencing a rapid evolution, heralding a new era of potential applications within the realm of professional content generation. The progression of LLMs in this domain necessitates the establishment of robust benchmarks to steer their development effectively. In this work, we introduce a novel benchmark dataset using unconstrained hallucination generation, comprising a dataset specifically curated for hallucinated news continuation, which encompasses in excess of 5,000 instances annotated at the keyword level. Additionally, we propose a secure, scalable, and user-friendly evaluation framework to facilitate comprehensive assessments. Through meticulous experimentation on eleven prominent LLMs, our study has unearthed a series of enlightening findings. Looking ahead, our research endeavors will persist in exploring the intricacies of hallucination phenomena within professional content generation. Concurrently, on the benchmarking front, we aspire to augment our datasets to encompass a more diverse spectrum of domains and linguistic variations, thereby broadening the applicability and relevance of our benchmarks.

### v2

@@ 3.6 Further Discussion
Each of the three evaluation forms possesses distinct advantages and drawbacks. Discriminative evaluation is often the method of choice for a range of standard benchmarks Li et al. (2023); Cheng et al. (2023). This approach is intuitive, and the construction of evaluation prompts is straightforward. Selective evaluation resembles discriminative evaluation but is marginally less demanding because it includes a reference option for contrast. In both discriminative and selective evaluations, certain models might be suspected of conjecturing answers from a few shots due to inadequate reasoning skills, which can undermine the reliability of the outcomes. Consequently, the use of explainable prompting becomes essential. Generative evaluation most closely mirrors real-world applications. However, the generated content is unrestricted, which poses challenges for even the most dependable reference-based evaluation techniques. Therefore, employing a combination of metrics simultaneously, including lexical evaluation based on token coverage and semantic evaluation based on textual similarity, is imperative.
@@
The foundational capabilities required of LLMs can be arrayed on a spectrum from simple to complex: generative, selective, and discriminative evaluation. Generative evaluation entails the direct invocation of parameters for continuation, bypassing the need for an extensive grasp of instructions. Selective evaluation necessitates a degree of inferential reasoning but offers comparative choices, rendering the level of difficulty moderate. Conversely, discriminative evaluation demands the precise retrieval of facts, thereby increasing the challenge.
4 Related Work
This section outlines benchmark datasets, their characteristics, and evaluation methodologies. These benchmarks are summarized in Table 5. For more related works, you may refer to Appendix C.
@@ 4 Related Work
Benchmark Generation Method: Base Dataset Annotation Metric Granularity Lang.
ChineseFactEval Wang et al. (2023a) Manual Manual Acc Sentence CN
CSK-PN Chen et al. (2023) Direct: Common KGs No Need Acc Word EN
FACTOR Muhlgay et al. (2023) CHG: Wiki, News Auto FACTOR Acc Sentence EN
FActScore Min et al. (2023) CHG: Wiki No Need FActScore by Human Short Sentence EN
@@
FactualityPrompts Lee et al. (2022) Direct: Wiki Auto NE Error, Entailment Document, Sentence EN
HADES Liu et al. (2022) CHG: Wiki Manual Acc, G-Mean, BSS, AUC, etc. Word EN
@@
HalluQA Cheng et al. (2023) CHG, Manual: TruthfulQA, Wiki Manual, Auto Non-hallucination Rate Sentence CN
HaLoCheck Elaraby et al. (2023) CHG No Need HaLoCheck, selfcheckGPT Sentence EN
HaluEval Li et al. (2023) CHG: Alpaca, HotpotQA, etc. Manual, Auto Acc Document EN
HILT Rawte et al. (2023) CHG: NYT, Politifact Manual HVI Word EN
@@
KoLA-KC Yu et al. (2024) Direct: Wiki, evolving dataset Auto BLEU, ROUGE Document EN
Med-HALT Pal et al. (2023) Direct: MedMCQA, PubMed, etc. No Need Acc, Pointwise Score All EN
PHD Yang et al. (2023b) CHG: Wiki Manual F1, Acc, Prec, Reca Document EN
SelfAware Yin et al. (2023) CHG: Quora, HowStuffWorks Manual F1, Acc Sentence EN
STSN Varshney et al. (2023) UHG Manual Acc, Prec, Reca Sentence, Concept EN
TruthfulQA Lin et al. (2022) Manual Manual Acc by Human or GPT-judge Sentence EN
@@
$\mathbb{UHG}$Eval (Ours) UHG: News Auto, Manual Acc, kwPrec, BERTScore, etc. Sentence, Keyword CN
@@
XSum Hallu Maynez et al. (2020) UHG: XSum Manual ROUGE, BERTScore, Acc, etc. Word, Document EN
@@
Table 5: Hallucination evaluation benchmarks sorted by name. In the Generation Method column, CHG refers to constrained hallucination generation, UHG refers to unconstrained hallucination generation, Manual indicates manually constructed, and Direct implies utilizing the base dataset without the need for generation. In the Annotation column, Auto denotes automatic machine annotation. In the Metric column, Acc, Prec, and Reca respectively indicate accuracy, precision, and recall. In the Lang. column, CN and EN respectively stand for Chinese and English.
4.1 Benchmark dataset construction
@@
Dataset construction usually involves three steps. Firstly, real-world texts for hallucination generation are collected, and most benchmarks directly use existing datasets, such as Wiki Muhlgay et al. (2023), Alpaca Li et al. (2023), PubMed Pal et al. (2023), etc. Secondly, hallucinations are generated usually by LLMs such as GPT3.5-Turbo, and most works use a constrained hallucination generation (CHG) paradigm. STSN Varshney et al. (2023) and XSum Hallu Maynez et al. (2020) are the only two benchmarks that use UHG as we do. Thirdly, it is not certain that the content generated by the LLMs actually contains hallucinations, and often requires annotation, which is mostly done by human involvement. There are also works using automatic machine labeling Muhlgay et al. (2023); Lee et al. (2022); Cheng et al. (2023). These are the basic methods for constructing datasets, but there are also some other paradigms, such as constructing the dataset purely using manual labor, e.g. ChineseFactEval Wang et al. (2023a), HADES Liu et al. (2022), TruthfulQA Lin et al. (2022), etc.
4.2 Benchmark dataset characteristics
@@
Regarding the granularity of hallucinations labeled in the datasets, most studies assess hallucinations at the sentence and document levels, while a few examine them at the word (or keyword, concept) level. Concerning language, most evaluation datasets are in English. To our knowledge, the only two Chinese benchmarks, ChineseFactEval Wang et al. (2023a) and HalluQA Cheng et al. (2023) contain only 125 and 450 questions, respectively. Given the notably limited size of these datasets, our work significantly enhances the pool of data available for Chinese hallucination evaluation.
4.3 Evaluation schemes
@@
Currently, building automatic metrics for evaluation is still dominant, and a small proportion of works use human evaluation Min et al. (2023); Lin et al. (2022); Maynez et al. (2020). In terms of specific evaluation metrics, most works adopt common classification metrics, e.g., F1, accuracy, precision, and recall. some other works construct their calculation methods, e.g., FACTOR Muhlgay et al. (2023), FActScore Min et al. (2023), HaLoCheck Elaraby et al. (2023), etc. However, the above metrics are rule-based and can only evaluate the ability of LLMs to classify hallucinations, but not the ability of LLMs to generate content without hallucinations. Thus, some benchmarks explore even further in generative evaluation. For example, KoLA Yu et al. (2024) evaluates knowledge creation (KC) using BLEU and ROUGE, and TruthfulQA Lin et al. (2022) evaluates hallucinations using a specially trained classifier, GPT-judge.
5 Conclusion
@@
LLMs are rapidly evolving, heralding a new era of potential applications within the realm of professional content generation. The progression of LLMs in this domain necessitates the establishment of robust benchmarks to steer their development effectively. In this work, we introduce a novel hallucination benchmark dataset using an unconstrained fashion, encompassing more than 5,000 instances annotated at the keyword level. Additionally, we propose a secure, scalable, and user-friendly evaluation framework to facilitate comprehensive assessments. Through meticulous experimentation on eleven prominent LLMs, our study has unearthed a series of enlightening findings. Looking ahead, our research endeavors will persist in exploring the intricacies of hallucination phenomena within professional content generation, aiming to further understand and enhance LLM capabilities.
Limitations
Dataset
Firstly, although we have utilized hallucination ranking, automatic labeling, human rechecking, and various other techniques mentioned in Appendix A to ensure the quality of data annotation, with over 5,000 data entries, there is still a possibility of labeling errors. We have mobilized the power of the open-source community to collectively improve our dataset. Additionally, the dataset creation process is flexible, allowing for dataset expansion into English and broader domains, such as mathematical reasoning and programming codes.
Framework
Although our framework simplifies the integration of LLMs through APIs or vLLM [^4] https://github.com/vllm-project/vllm, users seeking to utilize custom or diverse HuggingFace models may face initial hurdles. We need to further enhance the usability of our framework.
Constrained v.s. Unconstrained
We have determined that constrained generation cannot fully reflect real-world applications, but empirical analysis is required to prove this point. This may involve constructing a text classifier to determine the type of hallucination, followed by comparing the distribution of hallucinations in our dataset with those in other benchmark datasets to observe any significant deviations. We leave this for future work.
Appendix A The UHGEval Dataset
A.1 Dive into Human Rechecking Process
Least Hallucination Principle
The keyword-based labeling scheme has inherent limitations. Languages exhibit a dependency structure de Marneffe and Nivre (2019). For instance, in the phrase “The rainbow is black,” the words “rainbow” and “black” exhibit interdependence. One could contend that “black” is incorrect, while another could maintain that “rainbow” is erroneous, given that “night” is typically described as black. To address the challenges stemming from language dependency structures, we have adopted the Least Hallucination Principle. If a set of words can be selected, and their replacement with contextually appropriate words yields a reasonable sentence, then such a set of words is designated as a hallucinated word group. The words selected for annotation must meet the condition of comprising the minimal number of words in the group, as illustrated in Equation 1. In the equation, $\mathbf{W}$ is the set of keywords in a sentence, $\mathbf{w}$ is the hallucinated word group, $\text{correct}(\cdot)$ is the correction function that modifies hallucinated words to non-hallucinated words, and $\text{hallucinated}(\cdot)$ assesses whether a sentence composed of keywords hallucinated.
By this principle, within the phrase “Journey to the West is an American novel and one of the Four Great Classics,” the word “American” would be marked for annotation, as altering this single keyword to “Chinese” dispels the hallucination throughout the sentence.
Engagement of Annotators
Additionally, we acknowledge that hallucination annotation may become somewhat tedious. Consequently, annotators are integrated throughout the entire process, participating in discussions instead of solely evaluating the accuracy of machine annotations. This approach also yields benefits for our work. For example, an annotator with a journalism background offered valuable professional insights into pinpointing news-related hallucinations, emphasizing that fact increment is a critical aspect of news writing.
A.2 Analysis of Final Dataset
We developed a conversion rate chart to depict the transition from candidate hallucinations to the final dataset, as depicted in Fig. 6. The conversion rate can be interpreted as the likelihood of hallucinations occurring across various categories. Our observations indicate a higher likelihood of hallucinations in number-intensive and general news, whereas this likelihood is reduced in knowledge-intensive and document-intensive news.
Figure 6: Conversion rates from candidates to hallucinations.
By analyzing the hallucinated word cloud depicted in Fig. 7 for each news category, we can draw the following conclusions: Number-intensive news often includes numeric values that are challenging to remember, like 0.09% and 6:3, which pose difficulties for both LLMs and humans. General news encompasses a diverse vocabulary, featuring terms such as “social media” and “friendship,” which are often deemed less critical and thus challenging to incorporate into the training corpora of many LLMs. Knowledge-intensive news frequently features terms such as “according to incomplete statistics” and “key technology,” which are prevalent in technical literature. However, LLMs may not always use these terms appropriately. Document-intensive news often contains terms associated with official statements, such as “representation,” “president,” and “spokesperson.” This suggests that LLMs are susceptible to introducing unauthorized alterations to the content of documents.
Figure 7: Hallucinated keywords in different types of news
A.3 An example from the UHGEval dataset
Figure 8: An example from the UHGEval dataset. (In English)
Figure 9: An example from the UHGEval dataset. (In Chinese)
Appendix B Experiments
B.1 LLMs Employed in This Research
All LLMs used in this study are detailed in Table 6.
Model #Para. Publisher Date
GPT3.5-Turbo $*{}^{*}$ OpenAI $*{}^{*}$
GPT4-0613 NaN OpenAI 2023.06
ChatGLM2 6B Tsinghua 2023.06
Xinyu 7B IAAR&Xinhua 2023.06
InternLM 20B ShLab 2023.07
Baichuan2 13B Baichuan Inc. 2023.09
Baichuan2 53B Baichuan Inc. 2023.09
Qwen 14B Alibaba 2023.09
Aquila2 34B BAAI 2023.10
Xinyu2 70B IAAR&Xinhua 2023.10
GPT4-1106 NaN OpenAI 2023.11
Table 6: LLMs sorted by release date. All LLMs are chat models. Asterisk (*) denotes estimated value, NaN denotes no public data available, and 175B denotes 175billion.
GPT represents a series of LLMs developed by OpenAI OpenAI (2023). In this study, GPT3.5-Turbo, GPT4-0613, and GPT4-1106 are utilized. GLM constitutes a pre-training framework proposed by Tsinghua University Du et al. (2022), and the ChatGLM2-6B chat model is employed. InternLM serves as an open-source, lightweight training framework, with its development team releasing a spectrum of models utilizing this framework InternLM (2023); the InternLM-20B open-source chat model is utilized in the present work. Baichuan2 comprises a series of expansive, multilingual base language models Yang et al. (2023a), with both the open-source Baichuan2-7B chat model and the closed-source Baichuan2-53B chat model being employed in this investigation. Qwen encompasses a language model series characterized by distinct models with varying parameter counts Bai et al. (2023), and the Qwen-14B open-source chat model is utilized in the current study. Aquila2 represents a language model series devised by BAAI, noted for surpassing comparable models in terms of performance BAAI (2023), and the Aquila2-34B chat model is employed in this research.
Besides, the Xinyu series models are the results of a collaborative research and development effort between the Institute for Advanced Algorithms Research, Shanghai (IAAR, SH), and the State Key Laboratory of Media Convergence Production Technology and Systems of the Xinhua News Agency. Xinyu-7B is an augmented large-scale language model derived from the foundational model, BloomZ-7B Muennighoff et al. (2023) through continued pre-training, news-specific fine-tuning, and alignment optimization. And, Xinyu2-70B is developed based on the open-source LLaMA2-70B Touvron et al. (2023) framework, incorporating expansions to the Chinese lexicon, ongoing pre-training, and news-specific fine-tuning, thereby endowing it with a robust foundational capability in the news domain.
B.2 Evaluation Method
The evaluation of hallucinations can be decomposed into three principal dimensions: form, metric, and granularity. Form concerns how the model interacts with the evaluation dataset; metric refers to the precise computational approach utilized for performance assessment; and granularity signifies the depth of detail considered in the evaluation of hallucinations.
Form
This encompasses human evaluation, discriminative evaluation, selective evaluation, and generative evaluation, among others. Human evaluation entails the direct application of human judgment to determine if the model’s output contains hallucinations, representing a critical evaluation form Chang et al. (2024). However, the drawbacks of this approach are evident: evaluating too many data points is tantamount to annotating a new dataset, with the associated time and financial expenditures proving prohibitive.
Discriminative evaluation enables LLMs to respond with binary answers of “yes” or “no” Li et al. (2023); Cheng et al. (2023). Specifically, this evaluation modality involves presenting the LLM under scrutiny with an initial text followed by a continuation that may or may not include hallucinations. The LLM is tasked with producing a verdict as to the presence of hallucinations. Owing to the efficacy of few-shot prompting, this evaluation paradigm is relatively uncomplicated for LLMs to administer, as it facilitates the elicitation of the requisite responses. However, this method depends solely on the LLM’s ability to draw upon the knowledge encoded within its parameters, necessitating the concurrent application of knowledge and reasoning, and thus requiring a robust foundational model capacity.
Selective evaluation allows LLMs to tackle multiple-choice questions by choosing between option A or B, as exemplified by PandaLM Wang et al. (2024). Specifically, in selective evaluation, the LLM under evaluation is presented with an initial text followed by two continuations: one that includes hallucinations and another that does not. The LLM’s objective is to identify which of the two is hallucinated. This assessment method offers the LLM more contextual information than discriminative evaluation, thereby alleviating the burden of fact-checking and lessening the dependence on retrieving facts from its parameters. Consequently, this reduces the level of difficulty for the LLM.
However, both discriminative and selective evaluations encounter a substantial challenge. They are predicated on the assumption that “LLMs’s capacity to produce reliable text is contingent upon their discernment between hallucinated and non-hallucinated content.” These methods do not simulate the evaluation of the model’s output for hallucinations. Consequently, generative evaluation is crucial as it directly evaluates the presence of hallucinations in the text generated by the LLM under evaluation. However, the challenge arises from the fact that it is not feasible to automatically and accurately ascertain if the newly generated text is hallucinated; if it were, annotated datasets would be redundant. In scenarios of unrestrained text generation, this issue becomes increasingly complex. This complexity stems from the fact that text generated without constraints may introduce a multitude of entities and facts absent in the reference material, complicating the verification of their accuracy. Despite these hurdles, generative evaluation continues to be a predominant strategy in Natural Language Generation (NLG) tasks Novikova et al. (2017).
Metric
Metrics include classification metrics such as accuracy, precision, recall, and others, which are applicable to human evaluation, discriminative evaluation, and selective evaluation. Generative evaluation, on the other hand, encompasses both lexical and semantic metrics. Lexical metrics evaluate the extent of token overlap between the generated text and the reference information, including metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and the newly proposed metric by us, kwPrec. Semantic metrics gauge the similarity in meaning between sentences, with examples including BERTScore Zhang et al. (2020), GPT-judge Lin et al. (2022), and GPTScore Fu et al. (2023), among others.
Granularity
Evaluations can be conducted at both the sentence and keyword levels. Owing to our annotation methodology, our dataset is marked at the keyword level to signify instances of hallucinations. This approach affords a broader spectrum of possibilities for configuring the evaluation task, enabling the evaluated model to address the presence of hallucinations at either the keyword level, the sentence level, or even the document level.
B.3 UHGEval Framework in Detail
The framework comprises four ascending layers: the dependency layer, the evaluator layer, the core layer, and the interface layer.
The dependency layer defines the essential foundational components needed for the evaluation framework, including datasets, LLM hubs, and various metrics. Importantly, each component is designed for extensibility: datasets can be replaced with custom ones, LLMs can be integrated via APIs or platforms like Hugging Face [^5] https://huggingface.co/models, and metrics can be customized to fit specific needs.
The evaluator layer, constituting the second layer, centers on an abstract class, Evaluator, and its various implementations. Within this layer, three distinct types are implemented: GenerativeEvaluator, DiscriminativeEvaluator, and SelectiveEvaluator. Users may also engineer custom evaluators, contingent upon adherence to the interface specifications of the abstract class, necessitating merely three function overloads.
The core layer, representing the third stratum, comprises two principal modules: experiment.py and analyst.py. The former facilitates experiments involving multiple LLMs, evaluators, and processes, whereas the latter is tasked with the statistical analysis of experimental outcomes.
The interface layer, serving as the final layer, orchestrates the user’s interaction with $\mathbb{UHG}$Eval. To streamline the initiation process, a succinct 20-line demonstration is offered, alongside a run.py script for launching experiments through the command line.
Appendix C More Related Works
C.1 Large Language Models
Language models are pivotal in computer science, evolving from statistical language models to neural language models, to pre-trained language models (PLMs), and now to the current generation of LLMs. The advent of models such as ChatGPT has seen contemporary LLMs exhibit new capabilities in handling complex tasks. These models can manage few-shot tasks via in-context learning and tackle mixed tasks by following instructions Zhao et al. (2023).
LLMs can be classified according to two dimensions. The first dimension concerns the openness of the model weights. For example, open-source models include Meta’s LLaMA Touvron et al. (2023), Tsinghua University’s GLM Du et al. (2022), and Alibaba’s Qwen Bai et al. (2023), while closed-source models feature OpenAI’s GPT OpenAI (2023), Baidu’s ERNIE Bot Sun et al. (2021), and Anthropic’s Claude [^6] https://www.anthropic.com/index/introducing-claude, among others. The second dimension differentiates between the use of a PLM or a supervised fine-tuned (SFT) model for specific inferences. A PLM is a language model trained on extensive unlabeled textual data to discern underlying patterns, structures, and semantic knowledge within the corpus. Conversely, an SFT model involves further training a PLM with labeled datasets tailored to a specific task, to improve performance in that area. Many open-source models, including LLaMA, GLM, and Qwen, have made their PLM weights publicly available. For SFT models, users can access the chat variants of open-source models or the API services provided by closed-source models. In our research, we focus primarily on evaluating closed-source GPT series models and open-source Chinese chat models.
C.2 Hallucinations in LLM
Despite remarkable advancements in LLMs, they continue to encounter challenges, with hallucination being one of the most notable. Hallucination in language models refers to generating content that strays from factual accuracy, leading to unreliable outputs. Hallucinations occur when the generated content is not aligned with user input, deviates from the model’s previous outputs, or is at odds with established real-world knowledge Zhang et al. (2023).
Specific examples include inaccuracies in age, currency, scores, and other numerical values; citing fictional statements; inventing non-existent characters; and muddling timelines by merging events from different periods Rawte et al. (2023).
Regarding the causes of hallucinations, several factors can be responsible Zhang et al. (2023). One contributing factor is the use of inaccurate or incomplete training data. During training, LLMs fine-tune their parameters with vast quantities of text data. However, this data may be flawed, harboring errors, inaccuracies, or gaps in information. Another factor involves inconsistencies in contextual information. While LLMs typically consider previously generated context when producing content, challenges in managing long-term dependencies or understanding complex contexts can result in inconsistencies. Additionally, hallucinations can arise from lacking or erroneous world knowledge. Although LLMs gain considerable world knowledge via training data, they may be deficient in specific domain knowledge or misinterpret certain facts, leading to hallucinations. Furthermore, model limitations, including generation strategies and alignment methods, can also play a role in hallucinations during content creation.
Appendix D Figures in Chinese
Figure 10: Hallucinations from $\mathbb{UHG}$Eval. Using the IDs, you can locate the original news articles. (In English: Fig. 1)
Figure 11: The process of creating $\mathbb{UHG}$Eval. Steps 1 to 4 regarding the creation of the benchmark dataset are explained in Section 2; Step 5, concerning the evaluation framework, is detailed in Section 3. (In English: Fig. 2)
Figure 12: Labeling and rechecking. (In English: Fig. 4)
Appendix E Prompt Templates
In these templates, the orange text represents intent and instruction, the green text represents demonstrations, and the black text represents specific questions.
Figure 13: Prompt template for discriminative evaluation (keyword level)
Figure 14: Prompt template for discriminative evaluation (sentence level)
Figure 15: Prompt template for selective evaluation
Figure 16: Prompt template for generative evaluation